# Magpie: Alignment Data Synthesis from Scratch by Prompting Aligned LLMs with Nothing

### v1

@@ 3.2 Dataset Attributes
We use Llama-3-8B-Instruct to categorize the instances in Magpie-Pro (see Figure 7 in Appendix C.1 for detail). The prompts used to query Llama-3-8B-Instruct can be found in Appendix F. Our observations indicate that over half of the tasks in Magpie-Pro pertain to information seeking, making it the predominant category. This is followed by tasks involving creative writing, advice seeking, planning, and math. This distribution over the task categories aligns with the practical requests from human users [33].
Figure 4: The statistics of input difficulty and quality.
Attribute: Quality of Instructions.
We use the Llama-3-8B-Instruct model to assess the quality of each instruction in Magpie-Air and Magpie-Pro, categorizing them as ‘very poor’, ‘poor’, ‘average’, ‘good’, and ‘excellent’. We present the histograms of qualities for both datasets in Figure 4-(a). We have the following two observations. First, both datasets are of high quality, with the majority of instances rated ‘average’ or higher. In addition, the overall quality of Magpie-Pro surpasses that of Magpie-Air. We hypothesize that this is due to the enhanced capabilities of Llama-3-70B compared with Llama-3-8B.
Attribute: Difficulty of Instructions.
We use the Llama-3-8B-Instruct model to rate the difficulty of each instruction in Magpie-Air and Magpie-Pro. Each instruction can be labeled as ‘very easy’, ‘easy’, ‘medium’, ‘hard’, or ‘very hard’. Figure 4-(b) presents the histograms of the levels of difficulty for Magpie-Air and Magpie-Pro. We observe that the distributions across difficulty levels are similar for Magpie-Air and Magpie-Pro. Some instructions in Magpie-Pro are more challenging than those in Magpie-Air because Magpie-Pro is generated by a more capable model (Llama-3-70B-Instruct).
@@
Attribute: Instruction Similarity.
We quantify the similarity among instructions generated by Magpie to remove repetitive instructions. We measure the similarity using minimum neighbor distance in the embedding space. Specifically, we first represent all instructions in the embedding space using the all-mpnet-base-v2 embedding model. For any given instruction, we then calculate the minimum distance from the instruction to its nearest neighbors in the embedding space using Facebook AI Similarity Search (FAISS) [17]. The minimum neighbor distances of instructions in Magpie-Air after removing repetitions are summarized in Figure 5-(a).
Attribute: Quality of Responses.
We assess the quality of responses using a metric named reward difference. For each instance in our dataset, the reward difference is calculated as $r^{*}-r_{base}$, where $r^{*}$ is the reward assigned by a reward model to the response in our dataset, and $r_{base}$ is the reward assigned by the same model to the response generated by the Llama-3 base model for the same instruction. We use URIAL [35] to elicit responses from the base model. A positive reward difference indicates that the response from our dataset is of higher quality, and could potentially benefit instruction tuning. In our experiments, we follow [29] and use FsfairX-LLaMA3-RM-v0.1 [57] as our reward model. Our results on the reward difference are presented in Figure 5-(b).
@@ 3.3 Safety Analysis
We use Llama-Guard-2 [48] to analyze the safety of Magpie-Air and Magpie-Pro. Our results indicate that both datasets are predominantly safe, with less than 1% of the data potentially containing harmful instructions or responses. Please refer to Appendix C.2 for detailed safety analysis.
@@ 3.4 Cost Analysis
We perform experiments on a server with four NVIDIA A100-SXM4-80GB GPUs, an AMD EPYC 7763 64-Core Processor, and 512 GB of RAM, using the VLLM inference framework [28]. The models are loaded in the bfloat16 format.
When creating the 3M Magpie-Air dataset, our Magpie spent 1.55 and 50 hours to generate the instructions (Step 1) and responses (Step 2), respectively. For the 1M Magpie-Pro dataset, Magpie used 3.5 and 150 hours to generate the instructions (Step 1) and responses (Step 2), respectively. Compared to existing approaches to create instruction datasets, the pipeline of Magpie is fully automated without any human intervention or API access to advanced commercial models such as GPT-4 [1]. Consequently, Magpie is cost-effective and scalable. On average, implementing Magpie on a cloud server [^2] https://lambdalabs.com/service/gpu-cloud would incur costs of \$0.12 and \$1.1 per 1,000 data instances for Magpie-Air and Magpie-Pro, respectively.
@@ 3.5 Additional Analysis
Additional dataset analysis, including the impact of generation configurations on the quality and difficulty of the generated instructions, is detailed in Appendix C.3.
@@ 4 Performance Analysis
In this section, we evaluate the quality of datasets generated by Magpie by utilizing them to fine-tune model families including Llama-3 [40] and Qwen1.5 [3].
4.1 Experimental Setups.
Baselines for Instruction Tuning.
@@
We compare the family of datasets generated by Magpie with six state-of-the-art open-source instruction tuning datasets: ShareGPT [10], WildChat [64], Evol Instruct [58], UltraChat [16], OpenHermes [49], and Tulu V2 Mix [24]. ShareGPT and WildChat are representative human-written datasets containing 112K and 652K high-quality multi-round conversations between humans and GPT, respectively. Evol Instruct and UltraChat are representative open-source synthetic datasets. Following [39], we use the 208K sanitized version of Ultrachat provided by HuggingFace [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. OpenHermes and Tulu V2 Mix are crowd-sourced datasets consisting of a mix of diverse open-source alignment datasets, with 243K and 326K conversations, respectively. We note that to ensure fair comparison involving datasets of different sizes, we provide the results of Magpie-Pro-200K-Filtered and Magpie-Pro-100K-Filtered, which contains the first 200K and 100K conversations from Magpie-Pro-300K-Filtered. Detailed discussion on how to generate these datasets can be found in Appendix B.
Baselines for Instruction and Preference Tuning.
@@
We compare the models fine-tuned using data generated by Magpie with preference optimization baselines, including DPO [44], IPO [2], KTO [19] and ORPO [23]. Specifically, we follow [39] and use the models fine-tuned with the UltraChat dataset (for instruction tuning) and Ultrafeedback dataset (for preference optimization) [13].
Fine-Tuning Details.
@@
We follow [50] and use a cosine learning rate schedule with an initial learning rate of $2\times 10^{-5}$ when fine-tuning Llama-3 and Qwen1.5 models. The maximum sequence length is 8192. The fine-tuning process is conducted using four NVIDIA A100 GPUs with 80G memory, and the effective batch size is 32. The models are fine-tuned for 2 epochs. We follow the official instruction templates of each model.
Evaluation Benchmarks.
@@
We evaluate the performance of the fine-tuned models using two widely-adopted instruction-following benchmarks: AlpacaEval 2 [33] and Arena-Hard [32]. AlpacaEval 2 consists of 805 representative instructions chosen from real user interactions. Arena-Hard is an enhanced version of MT-Bench [66], containing 500 challenging user queries. Both benchmarks employ a GPT evaluator to assess responses generated by the model of interest and a baseline model. Specifically, we use GPT-4-Turbo (1106) and Llama-3-8B-Instruct as baselines for AlpacaEval 2. By default, Arena-Hard uses GPT-4 (0314) as its baseline model.
Metrics.
@@
We adopt two metrics to measure the capabilities of instruction-following of fine-tuned models. The first metric is the win rate (WR), which calculates the fraction of responses that are favored by the GPT evaluator. This metric is applied in both benchmarks including AlpacaEval 2 and Arena-Hard. The second metric is the length-controlled win rate (LC) [18], a debiased version of WR.
The GPT evaluator considers the lengths of responses generated by the baseline model and model under evaluation when computing LC. By accounting for response length, LC reduces its impact on the win rate. This metric is specifically applied to the AlpacaEval 2 benchmark [33].
@@
Detailed Experimental Setups. We provide more detailed descriptions of our experimental setups, including more fine-tuning details and decoding hyper-parameters in Appendix D.
@@ 4.2 Experimental Results
Table 2: This table compares the performance of models instruction-tuned on the Llama-8B base models using our datasets and baseline datasets. We observe that models fine-tuned with our datasets significantly outperform those fine-tuned with baseline datasets of the same order of magnitude in terms of data size. In addition, our fine-tuned models achieve comparable performance to the official aligned model, despite only undergoing SFT with a much smaller dataset. Numbers in bold indicate that Magpie outperforms the official Llama-3-8B-Instruct model.
@@
SFT +ShareGPT [10] 112K 9.73 7.2 0.81 27.26 18.32 1.18 6.5
+Evol Instruct [58] 143K 8.52 6.25 0.76 20.16 14.98 1.1 5.1
+OpenHermes [49] 243K 9.94 6.27 0.73 29.19 17.92 1.16 4.4
+Tulu V2 Mix [24] 326K 9.91 7.94 0.86 24.28 18.64 1.18 5.4
+WildChat[64] 652K 14.62 10.58 0.92 34.85 26.57 1.32 8.7
+UltraChat [16]
@@
+*PO   +UltraFeedback (DPO) [13, 44] 64K 18.36 17.33 1.14 44.42 42.36 1.46 14.8
  +UltraFeedback (IPO) [2, 13] 64K 17.46 16.13 1.11 41.66 38.45 1.43 14.2
  +UltraFeedback (KTO) [13, 19] 64K 15.81 14.62 1.05 41.33 38.32 1.42 12.2
  +UltraFeedback (ORPO) [13, 23] 64K 13.23 12.57 0.99 30.62 28.27 1.35 10.9
SFT +Magpie (Ours)
Air-300K-Raw 300K 21.99 21.65 1.21 48.63 48.06 1.42 15.8
Air-300K-Filtered 300K 22.66 23.99 1.24 49.27 50.8 1.44 14.9
Pro-300K-Raw 300K 21.65 22.19 1.2 49.65 50.84 1.42 15.9
Pro-100K-Filtered 100K 20.47 24.52 1.25 47.92 52.75 1.43 17.2
Pro-200K-Filtered 200K 22.11 26.02 1.26 51.17 56.76 1.41 15.9
Pro-300K-Filtered 300K 25.08 29.47 1.35 52.12 53.43 1.44 18.9
   Llama-3-8B-Instruct (SFT+RLHF) $>$10M [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct 22.92 22.57 1.26 50 50 - 20.6
Magpie datasets outperform others.
@@
In Table 2, we first compare the performance of Llama-3 models fine-tuned with datasets generated by Magpie against those fine-tuned with baseline datasets. Using the AlpacaEval 2 evaluation benchmark, we observe that both LC and WR of our fine-tuned models surpass those fine-tuned with baseline instruction datasets, regardless of the choice of the baseline model. This indicates that the datasets generated by Magpie are of higher quality, leading to significantly enhanced instruction-following capabilities. A similar observation is made when using the Arena-Hard evaluation benchmark. We highlight that the Llama-3 models fine-tuned with datasets generated by Magpie outperform even those models that have undergone preference optimization (e.g., instruction tuning combined with DPO), which emphasizes the high quality of data generated by Magpie.
@@
To investigate the advantages of Magpie across different task categories, we also compare the performance of models fine-tuned with Magpie-Pro compared with baseline datasets using WildBench benchmark [34]. This benchmark consists of 1024 tasks carefully selected from real-world human-LLM conversation logs. The results are demonstrated in Figure 6. We observe that Magpie consistently outperforms baseline datasets across categories.
@@
Models fine-tuned with data generated by Magpie achieve comparable performance to the official aligned model, but with fewer data. In Table 2, we compare the performance of models fine-tuned with data generated by Magpie against the official aligned model (Llama-3-8B-Instruct). We observe that the Llama-3-8B base model fine-tuned with data from Magpie outperforms Llama-3-8B-instruct using the AlpacaEval 2 benchmark. For example, using the Magpie-Pro-300K-Filtered dataset to fine-tune the Llama-3-8B base model results in WC 29.47% against GPT-4-Turbo (1106). Furthermore, when Llama-3-8B-Instruct is chosen as the baseline model of AlpacaEval 2, we observe that WC of Llama-3-8B base models fine-tuned with data from Magpie exceeds 50%, indicating a preference for our fine-tuned models over the official aligned model. Finally, we highlight that our fine-tuning process uses no more than 300K data, whereas the official aligned models are fine-tuned with more than 10M data samples. This demonstrates the high quality of the data generated by Magpie. Using the Arena-Hard benchmark, we observe that a 1.7% difference between the WR achieved using our fine-tuned model and the official aligned model. We attribute this discrepancy to the fraction of coding-related instructions in our dataset. We believe that this gap could be easily bridged as we increase the size of datasets.
Both data quantity and quality matter to capabilities of instruction-following. In what follows, we compare within the family of datasets generated by Magpie in Table 2. These datasets differ in sizes, deployment of filtering, and models used to generate data. We observe that as the size of dataset increases, the performance of fine-tuned model improves, indicating that data quantity plays a critical role in enhancing instruction-following capabilities. Furthermore, the model fine-tuned with Magpie-Pro-300K-Filtered outperform those fine-tuned with the same amount of raw data. This demonstrates the effectiveness of our filtering technique, and underscores the importance of data quality. Finally, we observe that the models fine-tuned with Magpie-Pro consistently outperform those fine-tuned with Magpie-Air. The reason is that Magpie-Pro is generated by the more capable model, i.e., Llama-3-70B-Instruct.
@@
Magpie can enhance the performance of other backbone models. Table 3 illustrates the efficacy of Magpie when applied to generate instruction dataset and fine-tune other backbone models, i.e., Qwen1.5-4B and Qwen1.5-7B. The results demonstrate that our fine-tuned models achieve better performance than the official aligned models, which have undergone instruction and preference tuning. These results underscore the effectiveness of Magpie and the quality of its generated instructions.
@@
Table 3: This table compares the performance of models instruction-tuned on the Qwen base models using the Magpie-Pro-300K-Filtered dataset and the official instruction-tuned models. The Qwen base model enhanced with Magpie consistently outperforms the official instruction-tuned model.
@@
Additional Experimental Results.
We defer additional experimental results and analysis of Magpie-Air-MT and Magpie-Pro-MT to Appendix E.1. Additionally, the performance of Magpie across various other benchmarks is reported in Appendix E.3.

### v2

@@ 3.2 Dataset Attributes
We use Llama-3-8B-Instruct to categorize the instances in Magpie-Pro (see Figure 9 in Appendix D.1 for detail). The prompts used to query Llama-3-8B-Instruct can be found in Appendix G. Our observations indicate that over half of the tasks in Magpie-Pro pertain to information seeking, making it the predominant category. This is followed by tasks involving creative writing, advice seeking, planning, and math. This distribution over the task categories aligns with the practical requests from human users (Li et al., 2023b).
Figure 3: The statistics of instruction difficulty and quality.
Attribute: Quality of Instructions. Similar to methods in (Chen et al., 2023), we prompt the Llama-3-8B-Instruct model to assess the quality of each instruction in Magpie-Air and Magpie-Pro, categorizing them as ‘very poor’, ‘poor’, ‘average’, ‘good’, and ‘excellent’. We present the histograms of qualities for both datasets in Figure 3-(a). We have the following two observations. First, both datasets are of high quality, with the majority of instances rated ‘average’ or higher. In addition, the overall quality of Magpie-Pro surpasses that of Magpie-Air. We hypothesize that this is due to the enhanced capabilities of Llama-3-70B compared with Llama-3-8B.
Attribute: Difficulty of Instructions. We use the Llama-3-8B-Instruct model to rate the difficulty of each instruction in Magpie-Air and Magpie-Pro. Each instruction can be labeled as ‘very easy’, ‘easy’, ‘medium’, ‘hard’, or ‘very hard’. Figure 3-(b) presents the histograms of the levels of difficulty for Magpie-Air and Magpie-Pro. We observe that the distributions across difficulty levels are similar for Magpie-Air and Magpie-Pro. Some instructions in Magpie-Pro are more challenging than those in Magpie-Air because Magpie-Pro is generated by a more capable model (Llama-3-70B-Instruct).
@@
Attribute: Instruction Similarity. We quantify the similarity among instructions generated by Magpie to remove repetitive instructions. We measure the similarity using minimum neighbor distance in the embedding space. Specifically, we first represent all instructions in the embedding space using the all-mpnet-base-v2 embedding model. For any given instruction, we then calculate the minimum distance from the instruction to its nearest neighbors in the embedding space using Facebook AI Similarity Search (FAISS) (Douze et al., 2024). The minimum neighbor distances of instructions in Magpie-Air after removing repetitions are summarized in Figure 4-(a).
Attribute: Quality of Responses. We assess the quality of responses using rewards assigned by a reward model, denoted as $r^{*}$. For each instance in our dataset, we also calculate reward difference as $r^{*}-r_{base}$, where $r_{base}$ is the reward assigned by the same reward model to the response generated by the Llama-3 base model for the same instruction. We use URIAL (Lin et al., 2023) to elicit responses from the base model. A positive reward difference indicates that the response from our dataset is of higher quality, and could potentially benefit instruction tuning. In our experiments, we follow Lambert et al. (2024) and choose FsfairX-LLaMA3-RM-v0.1 (Xiong et al., 2024) as the reward model. Our results on the reward difference are presented in Figure 4-(b).
@@ 3.3 Safety Analysis
We use Llama-Guard-2 (Team, 2024) to analyze the safety of Magpie-Air and Magpie-Pro. Our results indicate that both datasets are predominantly safe, with less than 1% of the data potentially containing harmful instructions or responses. Please see Appendix D.2 for detailed safety analysis.
@@ 3.4 Cost Analysis
We perform experiments on a server with four NVIDIA A100-SXM4-80GB GPUs, an AMD EPYC 7763 64-Core Processor, and 512 GB of RAM, using the VLLM inference framework (Kwon et al., 2023). The models are loaded in the bfloat16 format.
When creating the 3M Magpie-Air dataset, our Magpie spent 1.55 and 50 hours to generate the instructions (Step 1) and responses (Step 2), respectively. For the 1M Magpie-Pro dataset, Magpie used 3.5 and 150 hours to generate the instructions (Step 1) and responses (Step 2), respectively. Compared to existing approaches to create instruction datasets, the pipeline of Magpie is fully automated without any human intervention or API access to advanced commercial models such as GPT-4 (Achiam et al., 2023). Consequently, Magpie is cost-effective and scalable. On average, implementing Magpie on a cloud server [^2] https://lambdalabs.com/service/gpu-cloud would incur costs of \$0.12 and \$1.1 per 1,000 data instances for Magpie-Air and Magpie-Pro, respectively.
@@ 3.5 Additional Analysis
Additional dataset analysis, including the impact of generation configurations on the quality and difficulty of the generated instructions, is detailed in Appendix D.3. We provide ablation analysis on annotating models for assessing quality and difficulty in Appendix D.4.
@@ 4 Performance Analysis
In this section, we evaluate the quality of Magpie-generated datasets by utilizing them to align base models including Llama-3 (Meta, 2024), Qwen1.5 (Bai et al., 2023), and Qwen2 (Yang et al., 2024).
4.1 Experimental Setups
@@
Baselines for Supervised Fine-Tuning and Preference Optimization. We compare the family of instruction datasets generated by Magpie with eight SOTA open-source instruction datasets: ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), Evol Instruct (Xu et al., 2023a), UltraChat (Ding et al., 2023), GenQA (Chen et al., 2024), OpenHermes 1 (Teknium, 2023a), OpenHermes 2.5 (Teknium, 2023b), and Tulu V2 Mix (Ivison et al., 2023). ShareGPT and WildChat are representative human-written datasets containing 112K and 652K high-quality multi-round conversations between humans and GPT, respectively. Evol Instruct, UltraChat, and GenQA are representative open-source synthetic datasets. Following Meng et al. (2024), we use the 208K sanitized version of Ultrachat provided by HuggingFace [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. OpenHermes 1, OpenHermes 2.5, and Tulu V2 Mix are crowd-sourced datasets consisting of a mix of diverse open-source instruction datasets, with 243K, 1M, and 326K conversations, respectively. We also create an instruction dataset with 100K conversations using the Self-Instruct (Wang et al., 2023) and Llama-3-8B-Instruct model, denoted as Self-Instruct (Llama-3).
@@
We compare the models aligned using Magpie with preference optimization baselines using direct preference optimization (DPO) (Rafailov et al., 2023). Specifically, we follow Meng et al. (2024) and use the models fine-tuned with the UltraChat dataset (for instruction tuning) and Ultrafeedback dataset (for preference optimization) (Cui et al., 2023).
Magpie Setups. To demonstrate the quality of Magpie-generated instruction datasets for SFT, we select the first 300K Magpie-Air and Magpie-Pro raw datasets generated by Llama-3-8B-Instruct and Llama-3-70B-Instruct models, respectively. Apart from these raw datasets, we also applied the filters detailed in Appendix C and created two filtered datasets: Magpie-Air-Filtered and Magpie-Pro-Filtered, each contains 300K conversations. For preference optimization, we generate two additional datasets: Magpie-Air-DPO (generated by Llama-3-8B-Instruct) and Magpie-Pro-DPO (generated by Llama-3-70B-Instruct) with $k=5$ and $T=0.8$, each contains 100K conversations. We use RLHFlow/ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) as the reward model.
@@
Model Alignment Details. For supervised fine-tuning, we follow Touvron et al. (2023) and use a cosine learning rate schedule with an initial learning rate of $2\times 10^{-5}$ when fine-tuning Llama-3, Qwen1.5 and Qwen2 base models. The maximum sequence length is 8192. For DPO, we use a cosine learning rate of $5\times 10^{-7}$. The detailed parameters can be found in Appendix E.2. We follow the official instruction templates of each model.
@@
Evaluation Benchmarks. We evaluate the performance of the aligned models using two widely adopted instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b). AlpacaEval 2 consists of 805 representative instructions chosen from real user interactions. Arena-Hard is an enhanced version of MT-Bench (Zheng et al., 2023), containing 500 challenging user queries. Both benchmarks employ a GPT evaluator to assess responses generated by the model of interest and a baseline model. Specifically, we use GPT-4-Turbo (1106) and Llama-3-8B-Instruct as baselines for AlpacaEval 2. By default, Arena-Hard uses GPT-4 (0314) as its baseline model.
@@
Metrics. We adopt two metrics to measure the capabilities of instruction-following of fine-tuned models. The first metric is the win rate (WR), which calculates the fraction of responses that are favored by the GPT evaluator. This metric is applied in both benchmarks including AlpacaEval 2 and Arena-Hard. The second metric is the length-controlled win rate (LC) (Dubois et al., 2024), a debiased version of WR. The GPT evaluator considers the lengths of responses generated by the baseline model and model under evaluation when computing LC. By accounting for response length, LC reduces its impact on the win rate. This metric is specifically applied to the AlpacaEval 2 benchmark (Li et al., 2023b).
@@
More Experimental Setups. We provide more detailed descriptions of our experimental setups, including more model alignment details and benchmark decoding hyper-parameters in Appendix E.
@@ 4.2 Experimental Results
Table 1: This table compares the performance of models instruction-tuned on the Llama-3-8B base models using Magpie-generated datasets and baseline datasets. We observe that models aligned with our datasets significantly outperform those aligned with baseline datasets of the same order of magnitude in terms of data size. In addition, our fine-tuned models achieve comparable performance to the official aligned model, despite only undergoing SFT with a much smaller dataset. Numbers in bold indicate that Magpie outperforms the official Llama-3-8B-Instruct model.
@@
SFT +Self-Instruct (Llama-3) (Wang et al., 2023) 100K 7.21 5.18 0.7 17.86 12.73 1.05 4.0
+ShareGPT (Chiang et al., 2023) 112K 9.73 7.2 0.81 27.26 18.32 1.18 6.5
+Evol Instruct (Xu et al., 2023a) 143K 8.52 6.25 0.76 20.16 14.98 1.1 5.1
+OpenHermes 1 (Teknium, 2023a) 243K 9.94 6.27 0.73 29.19 17.92 1.16 4.4
+Tulu V2 Mix (Ivison et al., 2023) 326K 9.91 7.94 0.86 24.28 18.64 1.18 5.4
+WildChat (Zhao et al., 2024) 652K 14.62 10.58 0.92 34.85 26.57 1.32 8.7
+OpenHermes 2.5 (Teknium, 2023b) 1M 12.89 9.74 0.91 32.68 25.01 1.30 8.2
+GenQA (Chen et al., 2024) 6.47M 9.05 7.11 0.82 21.90 16.09 1.12 3.0
+UltraChat (Ding et al., 2023)
@@
+ DPO       +UltraFeedback((Cui et al., 2023)) 64K 18.36 17.33 1.14 44.42 42.36 1.46 14.8
SFT Magpie-Air-300K-Raw 300K 21.99 21.65 1.21 48.63 48.06 1.42 15.8
Magpie-Air-300K-Filtered
[…]
300K 22.66 23.99 1.24 49.27 50.8 1.44 14.9
+ DPO       +Magpie-Air-DPO 100K 45.48 50.43 1.48 75.06 79.64 1.18 35.9
SFT Magpie-Pro-300K-Raw 300K 21.65 22.19 1.2 49.65 50.84 1.42 15.9
Magpie-Pro-300K-Filtered
[…]
300K 25.08 29.47 1.35 52.12 53.43 1.44 18.9
+ DPO       +Magpie-Pro-DPO 100K 50.10 53.53 1.45 78.52 80.82 1.17 35.7
   Llama-3-8B-Instruct (SFT+DPO) $>$10M [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct 22.92 22.57 1.26 50 50 - 20.6
Magpie datasets outperform baselines with SFT only.
@@
In Table 1, we compare the performance of Llama-3 models fine-tuned with instruction datasets generated by Magpie against those supervised fine-tuned with baseline datasets. Using the AlpacaEval 2 benchmark, we observe that both the LC and WR of our supervised fine-tuned models surpass all those models fine-tuned with baseline SFT datasets. This indicates that the datasets generated by Magpie are of higher quality, leading to significantly enhanced instruction-following capabilities. A similar observation is made when using the Arena-Hard evaluation benchmark. We highlight that the Llama-3 base models supervised fine-tuned with instruction datasets generated by Magpie outperform even those models that have undergone preference optimization (i.e., STF followed by DPO), which further emphasizes the high quality of data generated by Magpie.
@@
To investigate the advantages of Magpie across different task categories, we also compare the performance of models fine-tuned with Magpie-Pro compared with baseline datasets using WildBench benchmark (Lin et al., 2024). This benchmark consists of 1024 tasks carefully selected from real-world human-LLM conversation logs. The results are demonstrated in Figure 5. We observe that Magpie consistently outperforms baseline datasets across categories.
@@
Models aligned with data generated by Magpie achieve comparable or even higher performance to the official aligned model, but with fewer data. In Table 1, we also compare the performance of models aligned with data generated by Magpie against the official aligned model (Llama-3-8B-Instruct). We observe that the Llama-3-8B base model supervised fine-tuned with data from Magpie outperforms Llama-3-8B-instruct using the AlpacaEval 2 benchmark. For example, when Llama-3-8B-Instruct is chosen as the baseline model of AlpacaEval 2, we observe that LC of Llama-3-8B base models fine-tuned with instruction data from Magpie exceeds 50%, indicating a preference for our SFT models over the official aligned model. In addition, when DPO is applied, our aligned model demonstrates remarkable performance gains. Specifically, it outperforms the official Llama-3-8B-Instruct model on both the AlpacaEval 2 and Arena-Hard benchmarks. Most notably, our model even surpasses GPT-4-Turbo(1106) on AlpacaEval 2. Finally, we highlight that our alignment process uses no more than 400K data, whereas the official aligned models are aligned with more than 10M data samples. This demonstrates the high quality of the data generated by Magpie.
@@
Magpie can enhance the performance of other backbone models. Table 2 illustrates the efficacy of Magpie when applied to generate instruction dataset and fine-tune other base models, i.e., Qwen2-1.5B, Qwen1.5-4B, and Qwen1.5-7B. The results demonstrate that our fine-tuned models achieve better performance than the official aligned models, which have undergone both supervised fine-tuning and preference tuning. These results underscore the effectiveness of Magpie and the quality of its generated instructions. In addition, we apply Magpie-generated datasets to align Llama-3.1-Minitron-4B-Width-Base (Sreenivas et al., 2024) and Llama-3.1-8B-Instruct (Dubey et al., 2024) using SFT followed by DPO. The resulting aligned model, which we term MagpieLM, achieves remarkable performance and ranks first among popular open-source instruction models with fewer than 10 billion parameters. The details of MagpieLM are deferred to Appendix B.
@@
Table 2: This table compares the performance of models instruction-tuned on the Qwen base models using the Magpie-Pro-300K-Filtered dataset and the official instruction-tuned models. The Qwen base model enhanced with Magpie outperforms the official instruction-tuned model.
@@
Qwen2-1.5B Qwen2-1.5B-Instruct 3.91 3.00 0.54 50 50 -
Base Model + Magpie 3.48 5.32 0.67 56.66 66.27 1.50
@@
Performance of Magpie on More Benchmarks. We report the performance of models supervised fine-tuned using Magpie-Air and Magpie-Pro, evaluated across a range of tasks featured on the Huggingface Open LLM Leaderboard (Beeching et al., 2023) in Table 3. The tasks includes MMLU (Hendrycks et al., 2020), ARC Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2021), WinoGrande (Levesque et al., 2012), and GSM8K (Cobbe et al., 2021). We also perform experiments on MMLU-Redux (Gema et al., 2024) with zero-shot prompting. Our experimental results demonstrate that models fine-tuned with Magpie-Air and Magpie-Pro achieve comparable performance to the official instruct model and other baselines.
We note that the performance of Magpie may degrade on reasoning tasks, which is attributed to the small proportion of reasoning instructions in Magpie-Air and Magpie-Pro datasets. In response, we provide a supplementary "booster" dataset containing 150K math, code, and reasoning instructions using the Magpie extension mentioned in Section 2.2. We combine this booster dataset with Magpie-Pro-300K-Filtered and create Magpie-Pro-Mix-Filtered. Experimental results presented in Table 3 demonstrate that the model supervised fine-tuned using the mixed dataset effectively addresses the initial weakness in reasoning tasks. Notably, this new model ranks among the top-3 of all model checkpoints, performing only slightly weaker than OpenHermes 2.5 (1M conversations) and Llama-3-8B-Instruct (>10M conversations). This significant improvement showcases the flexibility and adaptability of the Magpie framework in generating task-specific instruction data.
Table 3: This table compares the performance of models supervised-fine-tuned on Magpie-Air, Magpie-Pro, and Magpie-Pro-Mix against baselines and official instruct model across various downstream benchmarks. All models are supervised-fine-tuned on the Llama-8B base models.
Alignment Setup MMLU (5) ARC (25) HellaSwag (10) TruthfulQA (0) WinoGrande (5) GSM8K (5) MMLU-Redux (0) Average
ShareGPT 66.03 58.45 81.50 52.34 74.03 48.67 50.68 61.67
Evol Instruct 65.62 60.75 82.70 52.87 76.16 42.91 52.73 61.96
GenQA 63.45 58.53 79.65 48.85 74.03 43.14 51.87 59.93
OpenHermes 1 65.42 62.29 82.15 50.85 75.61 47.16 46.07 61.36
OpenHermes 2.5 65.70 61.86 82.53 51.35 76.09 67.02 46.07 66.24
Tulu V2 Mix 66.34 59.22 82.80 47.99 76.16 58.07 46.97 62.51
WildChat 65.95 59.22 81.39 53.18 75.30 48.75 52.59 62.34
UltraChat 65.23 62.12 81.68 52.76 75.53 50.57 50.75 62.66
Magpie-Air-300K-Filtered 64.45 61.01 79.90 53.48 72.38 52.24 52.34 62.25
Magpie-Pro-300K-Filtered 64.25 60.41 80.52 52.46 73.32 47.92 52.16 61.58
Magpie-Pro-Mix-Filtered 65.65 59.64 80.72 50.81 73.24 63.08 56.34 64.21
Llama-3-8B-Instruct 67.82 61.52 78.67 52.47 72.14 71.72 58.60 66.13
Additional Experimental Results. We defer additional experimental results and analysis of multi-turn datasets, i.e., Magpie-Air-MT and Magpie-Pro-MT, to Appendix F.1. We conduct a detailed comparison between Magpie and Self-Instruct in Appendix F.2. In addition, ablations on data quantity, quality, filter designs, and response generator are deferred in Appendices F.3, F.4, and F.5. Magpie model’s performance on trustworthiness benchmarks is reported in Appendices F.6.